OpenStudy (anonymous):

Translate into an inequality. DO NOT SOLVE. It took Alvin sixteen more minutes to drive to work today than it did yesterday. Tomorrow Alvin wants it to take 5 minutes less than yesterday so that his average drive time for the 3 days is no more than 30 minutes. How long did it take for Alvin to get to work each day?

OpenStudy (anonymous):

Let y = yesterday's time. Today = y + 16 Tomorrow = y - 5 [y + (y + 16) + (y - 5)] / 3 <= 30 Note: "<= 30" means less than or equal to 30

OpenStudy (anonymous):

Thank you!!! :)

OpenStudy (anonymous):

I don't know why my brain has a glitch between getting the problem into an equation. I look way too far into it I think.